# Occupancy-aware Trajectory Planning for Autonomous Valet Parking in Uncertain Dynamic Environments

### v1

@@ V-A Ablation Studies
We assess our different contributions by performing ablation studies on the spot occupancy estimator and strategy planner separately. In each case, we fix a specific module and vary components of the other, comparing simplified variants against our proposed design.
@@ V-A1 Spot Occupancy Estimator
  • Greedy approach assumes a horizon of $T=1$ in the belief estimation equations in Section III, thereby relying only on the instantaneous occupancy probabilities.
@@
(a) FoV when red vehicle moves out.
@@
Figure 4: Comparison of ego vehicle (green) trajectories and occupancy probabilities when a vehicle (red) vacates a spot with a nearby vacant spot (V). The Greedy and Identical estimators require 16.5 s to park, while our approach completes the maneuver in 11.5 s.
In Fig.4, we compare our estimator with the Greedy and Identical prediction method. The Greedy approach uses only the current occupancy, while ours predicts over a horizon by modeling vacant and occupied dynamics separately. Hence, our estimator significantly lowers the occupancy probability when a vehicle departs, unlike the Identical prediction (Fig. 4(e)). This yields a parking time of 11.5 s for our approach, compared to 16.5 s for others. Moreover, when the red vehicle briefly passes near the vacant spot V, its occupancy probability momentarily increases but later decreases since we incorporate velocity ((6) and (8)), whereas the probability for the initially occupied spot steadily decreases. Although spot V remains vacant, the planner avoids it since it is surrounded by static vehicles and selects a more accessible spot.
(a) FoV at initial time.
(b) Estimated beliefs of the vacant spot.
@@
Figure 5: Comparison of ego vehicle (green) trajectories and occupancy probabilities when a vehicle (red rectangle) passes a vacant and a pedestrian is entering another vehicle (red circle). The Position only method require 24 s to park, while our approach completes the maneuver in 11 s.
(a) FoV when red vehicle moves out.
@@
Figure 6: Comparison of ego vehicle (green) trajectories when there is no wait time in (18) and another vehicle (red) moves out of a spot. The No wait time strategy requires 27 s to park, while our strategy completes parking in 18 s.
@@
Figure 7: Comparison of trajectories when the ego vehicle (green) is initially outside the row of parking spots and a row is crowded with other dynamic agents (red). The lawn-mower strategy requires 43.5 s to park, while our strategy completes parking in 25.5 s.
In Fig.5, we compare our estimator with the Position only method as a vehicle passes a vacant spot. The Position only method assigns a higher occupancy probability to the spot (Fig.5(b)), incorrectly classifying the spot to be occupied from (17) ($P_{v}=0.3$). This misclassification leads the ego to skip the spot, resulting in longer parking times and less smooth trajectories (Fig.5(c)). The ego hence performs a backup maneuver before briefly waiting for a pedestrian crossing the road, who was initially outside the FoV (Fig.5(a)).
@@ V-A2 Strategy Planner
  • No wait time always sets $t_{w}=0$ in (18) so that it never pauses near a potentially vacant spot.
  • Lawn-mower performs pure exploration ($\mathcal{C}_{e}(\tau_{e})=-\mathcal{I}(\tau_{e})$} in (27)) when no promising spots are available as classified by (17).
Fig.6 compares the no wait time strategy with ours. Although the spot occupancy estimator predicts the initially occupied spot to be vacant in the future as illustrated in Fig. 4(d), the no wait strategy cannot plan a safe path and instead explores further before backing up, resulting in longer parking times. Our planner balances the trade-off between waiting and exploring, enabling the ego to park in an initially occupied spot (Fig. 6(c)) by permitting short, bounded waits. This behavior is especially valuable in dynamic environments where briefly waiting near a nearly vacant spot is more efficient than abandoning it.
In Fig.7, we show an example where the lawn-mower strategy wastes time exploring a crowded row despite feasible spots present in the adjacent row. Since it is purely exploration driven, the ego must wait for other vehicles both at the row entrance and inside, leading to larger parking time (Fig. 7(b)). Our strategy planner balances exploring the lot and exploiting smoother paths (Fig. 7(c)), thus avoiding costly detours and achieving shorter parking times.
TABLE III: Planning metrics for 50 randomized experiments across different methods: ablation studies in spot occupancy estimator and strategy planner, and existing work. The numbers in green denote the best value for each metric.
Method Runtime $\downarrow$ (s) Path length $\downarrow$ (m) Parking time $\downarrow$ (s) Average distance to closest dynamic obstacle
@@
Greedy $0.041\pm 0.002$ $34.603\pm 3.113$ $23.702\pm 5.211$ $3.335\pm 0.103$ $0.315\pm 0.041$ $4.655\pm 2.83$ $12.995\pm 2.341$ $0.061\pm 0.02$
Identical prediction $0.051\pm 0.002$ $32.589\pm 2.773$ $22.575\pm 4.097$ $3.687\pm 0.074$ $0.278\pm 0.092$ $4.302\pm 1.12$ $12.16\pm 1.154$ $0.056\pm 0.002$
Position only $0.049\pm 0.008$ $33.735\pm 5.298$ $21.539\pm 1.359$ $8.331\pm 0.949$ $0.314\pm 0.131$ $5.045\pm 2.45$ $14.6\pm 2.881$ $0.058\pm 0.012$
No wait time $0.065\pm 0.002$ $45.735\pm 5.328$ $28.524\pm 2.533$ $9.228\pm 1.25$ $0.264\pm 0.066$ $32.668\pm 4.987$ $18.989\pm 3.1$ $0.058\pm 0.015$
Lawn-mower $0.127\pm 0.03$ $41.858\pm 3.37$ $32.532\pm 3.58$ $3.273\pm 0.119$ $0.273\pm 0.013$ $12.536\pm 3.1$ $17.032\pm 1.725$ $0.059\pm 0.008$
InfPath - Traversal $0.059\pm 0.001$ $42.364\pm 3.15$ $26.725\pm 5.71$ $5.232\pm 0.579$ $0.261\pm 0.078$ $13.583\pm 1.541$ $16.86\pm 2.97$ $0.061\pm 0.034$
InfPath - MCBFT $0.281\pm 0.012$ $35.742\pm 2.05$ $23.35\pm 2.13$ $6.634\pm 1.02$ $0.276\pm 0.044$ $14.543\pm 2.146$ $14.655\pm 2.51$ $0.077\pm 0.009$
Rule-OBCA $0.416\pm 0.055$ $32.94\pm 2.11$ $25.195\pm 2.54$ $9.376\pm 0.974$ $0.335\pm 0.095$ $20.54\pm 1.981$ $15.375\pm 1.632$ $0.062\pm 0.04$
Ours $0.045\pm 0.005$ $24.001\pm 2.992$ $17.179\pm 4.333$ $8.564\pm 1.03$ $0.271\pm 0.052$ $3.513\pm 0.95$ $11.119\pm 2.2$ $0.024\pm 0.002$
@@ V-B Comparison with Existing Work
We benchmark our integrated spot occupancy estimator and strategy planner against three existing valet parking approaches: InfPath - Traversal [6], InfPath - MCBFT [6] and Rule-OBCA [8]. The InfPath method uses a Vanilla Bayes filter to estimate spot occupancies with a FoV observation model independent of distance from ego vehicle, and presents two planning methods: Traversal (Section IV-B of [6]) and Monte Carlo Bayes Filter Tree (MCBFT) (Section IV-C of [6]). While [6] is originally designed for exploration without dynamic obstacle avoidance, we apply our safety thresholds to discard unsafe paths and park at the closest available spot within the FoV.
  • InfPath - Traversal samples all feasible paths over a planning horizon of T = 5 secs, and selects the one that either parks at the closest spot or maximizes the information gain with obstacle avoidance. Information gain is computed using the occupancy estimation method presented in Section III of [6].
  • InfPath - MCBFT builds a policy tree from spot occupancy probabilities and selects paths using the Upper Confidence Bound.
  • Rule-OBCA chooses the closest available spot within the FoV and plans via rule-based dynamic collision avoidance with Model Predictive Control, using smooth nonlinear obstacle avoidance constraints [4].
Quantitatively, Table III shows that our approach substantially improves key planning metrics over existing approaches [6, 8] and variations of our approach (Section V-A), reducing path length and parking time while achieving lower smoothness costs and greater clearance from obstacles, resulting in more efficient and safer parking maneuvers.
@@ VI Conclusion and Future Work
We present a trajectory planning framework for autonomous valet parking that integrates a novel spot occupancy estimator and a strategy planner for uncertain, dynamic environments with limited FoV. The occupancy estimator distinguishes initially vacant and occupied spots to estimate future occupancy probabilities, while also incorporating motion of dynamic agents. This estimator is integrated with our strategy planner that balances information gain with goal-directed planning, and employs wait-and-go behaviors for promising spots. Simulation results show that our method outperforms existing approaches in efficiency, safety, and smoothness.
The following limitations in our work suggest directions for future work. First, our Hybrid A star planner limits control diversity. Hence, exploring sampling-based methods like MPPI [13] could enable richer strategies under tight constraints. Second, incorporating intent inference (e.g., parking, yielding, exiting) of other agents for interaction-aware planning could improve realism and safety. Third, integrating occlusion-aware sensing models could provide more realistic occupancy estimates. Finally, developing adaptive waiting behaviors by varying the maximum waiting time could lead to more efficient parking in dense parking lots.
APPENDIX
Uncertainty Propagation: We propagate the positional covariance at each timestep using a standard discrete-time model. The covariance matrix $\Sigma(k)$ is updated as
$\Sigma(k+1)=R(\theta(k))\left(\Sigma(k)+Q\right)R(\theta(k))^{\top},$
where $R(\theta(k))$ is the rotation matrix parameterized by the vehicle’s heading $\theta(k)$ and $Q$ is a diagonal matrix representing the process noise in the local frame. The initial covariance is $\Sigma(t)=\begin{bmatrix}\beta_{l}L&0\\ 0&\beta_{w}W\end{bmatrix}$, with $L$ and $W$ denoting the length and width of the vehicle, while $\beta_{l}$ and $\beta_{w}$ correspond to the respective scaling factors. If the dynamic agent is a pedestrian, then $\theta(k)=0\deg$, $\beta_{l}=\beta_{w}=1$, and we use a radius parameter to represent the size of the pedestrian instead of the length and width of the vehicle. To estimate the uncertainty in velocity, we approximate the instantaneous velocity as the finite difference $v^{j}(k)=\frac{y^{j}(k+1)-y^{j}(k)}{\Delta t}$, where $\Delta t$ is the time interval between successive time steps. The resulting covariance in velocity is

### v2

@@ V-A Ablation Studies
We assess our contributions by performing ablation studies on the spot occupancy estimator and strategy planner. In each case, we fix one module and vary components of the other, comparing the variants with our complete approach.
@@ V-A1 Spot Occupancy Estimator
  • Greedy approach assumes a horizon of $T=0.1~\textrm{s}$ (one step estimation) in Section III, primarily relying on the current occupancy probabilities.
@@
(a) FoV and predictions when red vehicle moves out.
@@
Figure 5: Comparison of ego vehicle (green) trajectories ((b) and (c)) and occupancy probabilities ((d) and (e)) when a vehicle (red) vacates a spot with a nearby vacant spot (V).
(a) FoV and predictions at initial time.
(b) Estimated beliefs of vacant spot.
@@
Figure 6: Comparison of ego vehicle (green) trajectories ((c) and (d)) and occupancy probabilities (b) when a vehicle (red rectangle) passes a vacant spot and a pedestrian (red circle) is entering another vehicle.
(a) FoV and predictions when red vehicle moves out.
@@
Figure 7: Comparison of ego vehicle (green) trajectories ((b) and (c)) when there is no wait time in (18) and another vehicle (red) moves out of a spot.
@@
Figure 8: Comparison of trajectories ((b) and (c)) when the ego vehicle (green) is initially outside the row of parking spots and a row is crowded with other dynamic agents (red).
Fig. 5 compares our estimator with the Greedy and Identical prediction methods. While Greedy relies only on current occupancy, our approach predicts over a horizon by modeling vacant and occupied dynamics separately, thus lowering occupancy probabilities when a vehicle departs (Fig. 5(e)) and reducing parking time to 11.5 s versus 16.5 s for others. When the red vehicle passes near the vacant spot V, its occupancy probability temporarily rises but later decreases due to velocity modeling in (6). Although spot V remains vacant, the planner avoids it due to surrounding static vehicles and chooses a more accessible spot.
Fig. 6 compares our estimator with the Position only method, which overestimates the occupancy probability (Fig. 6(b)) and causes the ego to skip the spot, leading to longer parking times of 24 s versus 11 s for our approach. The ego performs a backup maneuver (Fig. 6(c)) before briefly waiting for a pedestrian initially outside the FoV (Fig. 6(a)).
[…]
@@ V-A2 Strategy Planner
  • No wait time strategy always sets $t_{w}=0$ in (18) so that the ego never pauses near a potentially vacant spot.
  • Lawn-mower strategy inspired by [38] performs pure exploration ($\mathcal{C}_{e}(\tau_{e})=-\mathcal{I}(\tau_{e})$} in (27)) when no promising spots are available as classified by (17).
Fig. 7 compares the no wait time strategy with ours. Although our spot occupancy estimator predicts the initially occupied spot to become vacant (Fig. 5(d)), the no wait time strategy first explores further before backing up, resulting in a longer parking time of 27 s versus 18 s for our strategy. Our planner balances waiting and exploration, enabling the ego to park in an initially occupied spot (Fig. 7(c)) through short, bounded waits, which is effective in dynamic environments.
In Fig. 8, the Lawn-mower strategy wastes time exploring a crowded row. As a purely exploration-driven strategy, the ego waits for other vehicles, leading to a longer parking time of 43.5 s (Fig. 8(b)). Our strategy planner balances exploration and path efficiency, avoiding detours and reducing parking time to 25.5 s (Fig. 8(c)).
Table III shows that our complete approach (row 9) improves upon other strategy planner variants (rows 4–5), decreasing parking time and path length by at least 39%, while producing smoother paths, as indicated by lower heading rates and curvature.
TABLE III: Planning metrics for 50 randomized experiments across different methods: ablations on spot occupancy estimator and strategy planner, existing work and our approach. The numbers in green denote the best value for each metric.
No. Method Runtime $\downarrow$ (s) Path length $\downarrow$ (m) Parking time $\downarrow$ (s) Average distance to closest dynamic obstacle
@@
1 Greedy $0.041\pm 0.002$ $34.603\pm 3.113$ $23.702\pm 5.211$ $3.335\pm 0.103$ $0.315\pm 0.041$ $4.655\pm 2.83$ $12.995\pm 2.341$ $0.061\pm 0.02$
2 Identical prediction $0.051\pm 0.002$ $32.589\pm 2.773$ $22.575\pm 4.097$ $3.687\pm 0.074$ $0.278\pm 0.092$ $4.302\pm 1.12$ $12.16\pm 1.154$ $0.056\pm 0.002$
3 Position only $0.049\pm 0.008$ $33.735\pm 5.298$ $21.539\pm 1.359$ $8.331\pm 0.949$ $0.314\pm 0.131$ $5.045\pm 2.45$ $14.6\pm 2.881$ $0.058\pm 0.012$
4 No wait time $0.065\pm 0.002$ $45.735\pm 5.328$ $28.524\pm 2.533$ $9.228\pm 1.25$ $0.264\pm 0.066$ $32.668\pm 4.987$ $18.989\pm 3.1$ $0.058\pm 0.015$
5 Lawn-mower $0.127\pm 0.03$ $41.858\pm 3.37$ $32.532\pm 3.58$ $3.273\pm 0.119$ $0.273\pm 0.013$ $12.536\pm 3.1$ $17.032\pm 1.725$ $0.059\pm 0.008$
6 InfPath - Traversal $0.059\pm 0.001$ $42.364\pm 3.15$ $26.725\pm 5.71$ $5.232\pm 0.579$ $0.261\pm 0.078$ $13.583\pm 1.541$ $16.86\pm 2.97$ $0.061\pm 0.034$
7 InfPath - MCBFT $0.281\pm 0.012$ $35.742\pm 2.05$ $23.35\pm 2.13$ $6.634\pm 1.02$ $0.276\pm 0.044$ $14.543\pm 2.146$ $14.655\pm 2.51$ $0.077\pm 0.009$
8 Rule-OBCA $0.416\pm 0.055$ $32.94\pm 2.11$ $25.195\pm 2.54$ $9.376\pm 0.974$ $0.335\pm 0.095$ $20.54\pm 1.981$ $15.375\pm 1.632$ $0.062\pm 0.04$
9 Ours $0.045\pm 0.005$ $24.001\pm 2.992$ $17.179\pm 4.333$ $8.564\pm 1.03$ $0.271\pm 0.052$ $3.513\pm 0.95$ $11.119\pm 2.2$ $0.024\pm 0.002$
@@ V-B Comparison with Existing Work
We benchmark our valet parking framework—combining a spot occupancy estimator and strategy planner—against three existing approaches: InfPath - Traversal [23], InfPath - MCBFT [23], and Rule-OBCA [20]. InfPath estimates spot occupancies using a vanilla Bayes filter that ignores dynamic agents and employs an observation model independent of distance from ego vehicle, and presents two planning strategies: Traversal (Section IV-B of [23]) and Monte Carlo Bayes Filter Tree (MCBFT, Section IV-C of [23]). Although [23] focuses on exploration without dynamic obstacle avoidance, we enforce safety thresholds to discard unsafe paths and select the closest vacant spot according to (17).
  • InfPath - Traversal samples feasible trajectories over a planning horizon of $T=5$ s, and selects trajectories that either parks at the closest spot or explores using information gain that is computed using the occupancy estimation method presented in Section III of [23].
  • InfPath - MCBFT builds a policy tree from spot occupancy probabilities and selects trajectories using the Upper Confidence Bound.
  • Rule-OBCA selects the nearest available spot within the FoV and executes rule-based dynamic collision avoidance using Model Predictive Control with smooth nonlinear constraints [7].
Quantitatively, rows 6–9 in Table III show that our approach outperforms existing methods [23, 20], reducing path length and parking time by at least 25% while achieving lower smoothness costs and maintaining safe clearance from obstacle. The runtime presented in Table III is for our combined spot occupancy estimator and strategy planner.
@@ VI Conclusion and Future Work
We present a trajectory planning framework for autonomous valet parking with limited FoV that integrates a novel spot occupancy estimator and a strategy planner in uncertain, dynamic environments. We model initially vacant and occupied spots separately to estimate future occupancy probabilities, while also incorporating the motion of dynamic agents. The estimator is integrated with our strategy planner that balances information gain with goal-directed planning, and employs wait-and-go behaviors for promising spots. Simulation results show that our method outperforms existing approaches in efficiency, safety, and smoothness.
The following limitations in our work suggest directions for future research. First, the Hybrid A planner limits steering and velocity samples. Thus, exploring sampling-based methods like MPPI [32] could enable richer control strategies. Second, incorporating intent inference (e.g., parking, yielding, exiting) for other vehicles and evaluating the performance across a wide range of learned trajectory prediction models could enhance interaction-aware planning, thereby improving realism and safety. Third, occlusion-aware sensing models could provide more realistic occupancy estimates. Additionally, developing adaptive waiting behaviors based on lot density may improve efficiency. Finally, implementing our approach in high-fidelity simulators with interactive agents and real-world environments is our future goal.